### v1

# Finite-Time Analysis of Simultaneous Double Q-learning

Abstract
$Q$-learning is one of the most fundamental reinforcement learning (RL) algorithms. Despite its widespread success in various applications, it is prone to overestimation bias in the $Q$-learning update. To address this issue, double $Q$-learning employs two independent $Q$-estimators which are randomly selected and updated during the learning process. This paper proposes a modified double $Q$-learning, called simultaneous double $Q$-learning (SDQ), with its finite-time analysis. SDQ eliminates the need for random selection between the two $Q$-estimators, and this modification allows us to analyze double $Q$-learning through the lens of a novel switching system framework facilitating efficient finite-time analysis. Empirical studies demonstrate that SDQ converges faster than double $Q$-learning while retaining the ability to mitigate the maximization bias. Finally, we derive a finite-time expected error bound for SDQ.
1 Introduction
Reinforcement learning (RL) is a class of learning algorithms for finding an optimal policy in unknown environments through experiences [1]. $Q$-learning [2] is one of the most popular RL algorithms, which finds an optimal policy by estimating the optimal $Q$-function. Its effectiveness has been established through many successful applications [3, 4, 5, 6] as well as theoretical convergence analysis [7, 8, 9, 10, 11, 12, 13].
Despite its empirical successes and theoretical achievements, $Q$-learning is known to suffer from overestimation in the $Q$-estimator, known as the maximization bias [1], which slows down the convergence of the corresponding policy to an optimal policy especially in environments with a large number of actions and heterogeneous action spaces. To overcome this obstacle, the so-called double $Q$-learning was proposed in [14], which empirically proved that the maximization bias can be reduced by using double $Q$-estimators instead of the single $Q$-estimator. Since its introduction, double $Q$-learning has been successfully applied in practice [15, 16, 17], and analyzed thoroughly in [18, 19]. However, from a practical standpoint, double $Q$-learning introduces an additional random selection step between the two $Q$-estimators, which can slow down its convergence speed compared to standard $Q$-learning despite its advantage in mitigating maximization bias.
Motivated by the aforementioned discussion, this paper proposes a modified double $Q$-learning called simultaneous double $Q$-learning (SDQ), which departs from the original in two key aspects: 1) Elimination of random selection: It dispenses with the need for random selection between the two $Q$-estimators, a step that can slow down convergence in the original double $Q$-learning. Through experiments, effectiveness of SDQ is demonstrated in this paper; 2) Different roles of $Q$-estimators: The role of each $Q$-estimator is distinct from those employed in the traditional double $Q$-learning algorithm in the sense that the roles of the two $Q$-estimators are swapped in the update rule. This change allows us to view SDQ through the lens of switching system models [11, 12, 13], enabling more efficient finite-time analysis. Moreover, we provide expected finite-time error bounds of SDQ instead of error bounds with high probability constraints [18, 19] in the literature. To our knowledge, this is the first such analysis for double $Q$-learning.
To establish the finite-time error bounds, a novel analysis framework is developed in this paper. In particular, SDQ is modelled as a switching system [11, 12, 13], which captures the dynamics of double $Q$-learning as a discrete-time switching system model. For finite-time convergence analysis, two comparison systems – termed the lower comparison system and the upper comparison system – are derived to bound the behavior of the original switching system. Through convergence of these comparison systems, the following expected error bound is derived:
$\text{max}\{\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}],\,\mathbb{E}[\lVert Q_{k}^{B}-Q^{*}\rVert_{\infty}]\}$ $\leq\frac{120\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\rho^{k-4}k^{4}\lvert S\times A\rvert^{3/2}}{(1-\gamma)},$
where $\lvert S\times A\rvert$ is the number of the state-action pairs, $\gamma$ is the discount factor, $d_{\text{min}}$ is the minimum state-action occupation frequency, $\alpha\in(0,1)$ is the constant step-size, $Q^{*}$ is a vectorized optimal $Q$-function, $Q^{A}_{k}$ and $Q^{B}_{k}$ are its double estimates at the current time $k$, and $\rho\coloneqq 1-\alpha d_{\text{min}}(1-\gamma)\in(0,1)$ is the exponential decay rate.
Although the switching system model has been first introduced in [11, 12, 13], we extend this view to double $Q$-learning and provide a new finite-time analysis. We note that this extension is not trivial due to more complex structures of double $Q$-learning, which includes an extra error term between the two $Q$-estimators that makes the analysis much more challenging compared to the standard $Q$-learning. Therefore, the techniques used in the previous studies cannot be directly applied to double $Q$-learning. In this paper, new approaches have been developed to overcome this challenge. Details on the proposed analysis can be found in Section 5, 6. Finally, the main contributions are summarized as follows:
  1. (a)
    Simultaneous double $Q$-learning (SDQ) is proposed to address maximization bias while exhibiting favorable convergence properties. Moreover, this modification enables viewing double $Q$-learning through the lens of a switching system, enabling more efficient finite-time analysis;
  2. (b)
    Based on the switching system model, novel finite-time analysis techniques and new expected error bounds are proposed for the SDQ. Moreover, the analysis frameworks introduced in this paper provide new theoretical perspectives and additional insights on double $Q$-learning and related algorithms.
2 Related Works
There have been relatively little researches on the convergence analysis of double $Q$-learning. The first convergence proof of double $Q$-learning has been provided by [14], but it only established asymptotic convergence. Finite-time convergence results have been recently given in [18, 19]. Both conducted convergence analysis on synchronous and asynchronous double $Q$-learning with non-i.i.d. (independent and identically distributed) observation model. In this paper, we analyze the convergence properties of a variant of double $Q$-learning where the $Q$-estimators serve different roles compared to the original approach. We specifically focus on asynchronous double $Q$-learning with modifications and i.i.d. observation model for our analysis. Regarding the step-size, the previous analysis in [18, 19] imposes more restrictive ranges of step-sizes for their analysis. However, the proposed analysis allows more flexible step-sizes $\alpha\in(0,1)$. Moreover, the existing studies have derived error bounds with high probability constraints. However, the proposed results are presented in terms of the expected error bounds, which provide additional insights and complement the existing analysis.
3 Preliminaries
3.1 Markov Decision Problem
We consider the infinite-horizon discounted Markov decision problem (MDP), where the agent sequentially takes actions to maximize cumulative discounted rewards. In a Markov decision process with the state-space ${\cal S}:=\{1,2,\ldots,|{\cal S}|\}$ and action-space ${\cal A}:=\{1,2,\ldots,|{\cal A}|\}$, the decision maker selects an action $a\in{\cal A}$ with the current state $s$, then the state transits to a state $s^{\prime}$ with probability $P(s,a,s^{\prime})$, and the transition incurs a reward $r(s,a,s^{\prime})$, where $P(s,a,s^{\prime})$ is the state transition probability from the current state $s\in{\cal S}$ to the next state $s^{\prime}\in{\cal S}$ under action $a\in{\cal A}$, and $r(s,a,s^{\prime})$ is the reward function. For convenience, we consider a deterministic reward function and simply write $r(s_{k},a_{k},s_{k+1})=:r_{k},k\in\{0,1,\ldots\}$.
A deterministic policy, $\pi:{\cal S}\to{\cal A}$, maps a state $s\in{\cal S}$ to an action $\pi(s)\in{\cal A}$. The objective of the MDP is to find a deterministic optimal policy, $\pi^{*}$, such that the cumulative discounted rewards over infinite time horizons is maximized, i.e.,
$\pi^{*}:=\operatorname*{arg\,max}_{\pi\in\Theta}{\mathbb{E}}\left[\left.\sum_{k=0}^{\infty}{\gamma^{k}r_{k}}\right|\pi\right],$
where $\gamma\in[0,1)$ is the discount factor, $\Theta$ is the set of all admissible deterministic policies, $(s_{0},a_{0},s_{1},a_{1},\ldots)$ is a state-action trajectory generated by the Markov chain under policy $\pi$, and ${\mathbb{E}}[\cdot|\pi]$ is an expectation conditioned on the policy $\pi$. The $Q$-function under policy $\pi$ is defined as
$Q^{\pi}(s,a)={\mathbb{E}}\left[\left.\sum_{k=0}^{\infty}{\gamma^{k}r_{k}}\right|s_{0}=s,a_{0}=a,\pi\right],\quad s\in{\cal S},a\in{\cal A},$
and the optimal $Q$-function is defined as $Q^{*}(s,a)=Q^{\pi^{*}}(s,a)$ for all $s\in{\cal S},a\in{\cal A}$. Once $Q^{*}$ is known, then an optimal policy can be retrieved by the greedy policy $\pi^{*}(s)=\operatorname*{arg\,max}_{a\in{\cal A}}Q^{*}(s,a)$. Throughout, we assume that the MDP is ergodic so that the stationary state distribution exists and the MDP is well posed.
3.2 Switching system
Since the switching system [20, 21] is a special form of nonlinear systems [22], we first consider the nonlinear system
$x_{k+1}=f(x_{k}),\quad x_{0}=z\in{\mathbb{R}}^{n},\quad k\in\{1,2,\ldots\},$ (1)
where $x_{k}\in{\mathbb{R}}^{n}$ is the state and $f:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ is a nonlinear mapping. An important concept in dealing with the nonlinear system is the equilibrium point. A point $x=x^{*}$ in the state-space is said to be an equilibrium point of (1) if it has the property that whenever the state of the system starts at $x^{*}$, it will remain at $x^{*}$ [22]. For (1), the equilibrium points are the real roots of the equation $f(x)=x$. The equilibrium point $x^{*}$ is said to be globally asymptotically stable if for any initial state $x_{0}\in{\mathbb{R}}^{n}$, $x_{k}\to x^{*}$ as $k\to\infty$.
Next, let us consider the particular system, called the linear switching system,
$x_{k+1}=A_{\sigma_{k}}x_{k},\quad x_{0}=z\in{\mathbb{R}}^{n},\quad k\in\{0,1,\ldots\},$ (2)
where $x_{k}\in{\mathbb{R}}^{n}$ is the state, $\sigma\in{\mathcal{M}}:=\{1,2,\ldots,M\}$ is called the mode, $\sigma_{k}\in{\mathcal{M}}$ is called the switching signal, and $\{A_{\sigma},\sigma\in{\mathcal{M}}\}$ are called the subsystem matrices. The switching signal can be either arbitrary or controlled by the user under a certain switching policy. Especially, a state-feedback switching policy is denoted by $\sigma_{k}=\sigma(x_{k})$. A more general class of systems is the affine switching system
$x_{k+1}=A_{\sigma_{k}}x_{k}+b_{\sigma_{k}},\quad x_{0}=z\in{\mathbb{R}}^{n},\quad k\in\{0,1,\ldots\},$
where $b_{\sigma_{k}}\in{\mathbb{R}}^{n}$ is the additional input vector, which also switches according to $\sigma_{k}$. Due to the additional input $b_{\sigma_{k}}$, its stabilization becomes much more challenging.
3.3 Double Q-learning
Double $Q$-learning [14] is a variant of $Q$-learning [2], which can overcome the maximization bias in its update by updating one of the two $Q$-estimators $Q^{A}_{k}$ and $Q^{B}_{k}$, which is selected randomly. Therefore, the corresponding update can be presented as follows:
$Q_{k+1}^{A}(s_{k},a_{k})$ $=\zeta_{k}Q_{k}^{A}(s_{k},a_{k})+\alpha_{k}\zeta_{k}\{r_{k}+\gamma Q_{k}^{B}(s_{k+1},\text{argmax}_{a\in A}Q_{k}^{A}(s_{k},a))-Q_{k}^{A}(s_{k},a_{k})\},$
$Q_{k+1}^{B}(s_{k},a_{k})$ $=(1-\zeta_{k})Q_{k}^{B}(s_{k},a_{k})+\alpha_{k}(1-\zeta_{k})\{r_{k}+\gamma Q_{k}^{A}(s_{k+1},\text{argmax}_{a\in B}Q_{k}^{B}(s_{k},a))$
$\quad-Q_{k}^{B}(s_{k},a_{k})\},$ (3)
where the updates for $Q^{A}_{k}$ and $Q^{B}_{k}$ are chosen randomly with the Bernoulli random variable $\zeta_{k}$ which satisfies $\mathbb{P}(\zeta_{k}=0)=\mathbb{P}(\zeta_{k}=1)=0.5$. By eliminating the max operator in its updates, it is known to reduce effectively the maximization bias.
3.4 Assumption and Definition
Throughout, we make the following standard assumptions, which are widely adopted in the reinforcement learning (RL) literature. We consider the scenario where the data is collected by the RL agent following a behavior policy $\beta$, which is the policy that the agent actually uses to interact with the environment and collect experiences. Moreover, we assume that the state-action pair at each time step is sampled i.i.d. from the stationary state distribution $p$ and the behavior policy $\beta$, respectively. Under these assumptions, the joint state-action distribution can be represented by
$d(s,a)=p(s)\beta(a|s),\quad(s,a)\in S\times A.$
Moreover, the following assumptions will be adopted.
Assumption 3.1.
$d(s,a)>0$ holds for all $s\in S,a\in A$.
Assumption 3.2.
The step-size is a constant $\alpha\in(0,1).$
Assumption 3.3.
We have
$\max_{(s,a,s^{\prime})\in S\times A\times S^{\prime}}\lvert r(s,a,s^{\prime})\rvert=R_{\text{max}}\leq 1.$
Assumption 3.4.
The initial iterate $Q_{0}$ satisfies $\lVert Q_{0}\rVert_{\infty}\leq 1.$
3.1 is required for sufficient explorations, and 3.3 and 3.4 are introduced without loss of generality and for simplicity of the analysis. The following quantities will be frequently used in this paper; hence, the following notations are defined for convenience.
Definition 3.5.
  1) Maximum state-action occupancy frequency:
$d_{\text{max}}\coloneq\max_{(s,a)\in S\times A}d(s,a)\in(0,1).$
2) Minimum state-action occupancy frequency:
$d_{\text{min}}\coloneq\min_{(s,a)\in S\times A}d(s,a)\in(0,1).$
3) Exponential decay rate:
$\rho\coloneq 1-\alpha d_{\text{min}}(1-\gamma).$ (4)
Under 3.2, the decay rate satisfies $\rho\in(0,1)$. Throughout the paper, we will use the following compact notations for dynamical system representations:
$P=\begin{bmatrix}P_{1}\\ \vdots\\ P_{\lvert A\rvert}\end{bmatrix},R=\begin{bmatrix}R_{1}\\ \vdots\\ R_{\lvert A\rvert}\end{bmatrix},Q=\begin{bmatrix}Q(\cdot,1)\\ \vdots\\ Q(\cdot,\lvert A\rvert)\end{bmatrix},$
$D_{a}=\begin{bmatrix}d(1,a)&&\\ &\dots&\\ &&d(\lvert S\rvert,a)\end{bmatrix},\quad D=\begin{bmatrix}D_{1}&&\\ &\dots&\\ &&D_{\lvert A\rvert}\\ \end{bmatrix},$
where $P_{a}=P(\cdot\lvert a,\cdot)\in\mathbb{R}^{\lvert S\rvert\times\lvert S\rvert}$, $Q(\cdot,a)\in\mathbb{R}^{\lvert S\rvert}$, $a\in\textit{A}$, and $R_{a}(s)\coloneqq\mathbb{E}\left[r(s,a,s^{\prime})\lvert s,a\right]$. Note that $P\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert\times\lvert S\rvert}$, $R\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert}$, $Q\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert}$, and $D\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert\times\lvert S\rvert\lvert A\rvert}$. Using these notations, $Q$-function is encoded as a single vector $Q\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert}$, which enumerates $Q(s,a)$ for all $s\in S$ and $a\in A$. The single value $Q(s,a)$ can be written as $Q(s,a)=(e_{a}\otimes e_{s})^{T}Q$, where $e_{s}\in\mathbb{R}^{\lvert S\rvert}$ and $e_{a}\in\mathbb{R}^{\lvert A\rvert}$ are $s$-th basis vectors (all components are 00 except for the $s$-th component which is 1) and $a$-th basis vectors, respectively, and $\otimes$ represents the Kronecker product. Under 3.2, $D$ is a non-singular diagonal matrix with strictly positive diagonal elements. For any stochastic policy, $\pi:S\rightarrow\Delta_{\lvert A\rvert}$, where $\Delta_{\lvert A\rvert}$ is the set of all probability distributions over A, let us define the matrix
$\Pi^{\pi}\coloneqq\begin{bmatrix}\pi(1)^{T}\otimes e_{1}^{T}\\ \pi(2)^{T}\otimes e_{2}^{T}\\ \vdots\\ \pi(\lvert S\rvert)^{T}\otimes e_{\lvert S\rvert}^{T}\\ \end{bmatrix}\in\mathbb{R}^{\lvert S\rvert\times\lvert S\rvert\lvert A\rvert,}$ (5)
where $e_{s}\in\mathbb{R}^{\lvert S\rvert}$. Then, it is well known that $P\Pi^{\pi}\in{\mathbb{R}}^{|{\cal S}||{\cal A}|\times|{\cal S}||{\cal A}|}$ is the transition probability matrix of the state-action pair under policy $\pi$. If we consider a deterministic policy, $\pi:{\cal S}\to{\cal A}$, the stochastic policy can be replaced with the corresponding one-hot encoding vector $\vec{\pi}(s):=e_{\pi(s)}\in\Delta_{|{\cal A}|},$ where $e_{a}\in{\mathbb{R}}^{|{\cal A}|}$, and the corresponding action transition matrix is identical to (5) with $\pi$ replaced with $\vec{\pi}$. For any given $Q\in{\mathbb{R}}^{|{\cal S}||{\cal A}|}$, denote the greedy policy w.r.t. $Q$ as $\pi_{Q}(s):=\operatorname*{arg\,max}_{a\in{\cal A}}Q(s,a)\in{\cal A}$. We will use the following shorthand frequently: $\Pi_{Q}:=\Pi^{\pi_{Q}}$.
The boundedness of $Q$-learning iterates [23] plays an important role in our analysis.
Lemma 1.
[23] If the step-size is less than one, then for all $k\geq 0$
$\lVert Q_{k}\rVert_{\infty}\leq Q_{\text{max}}=\frac{\text{max}{\{R_{\text{max}},\text{max}_{(s,a)\in S\times A}Q_{0}(s,a)\}}}{1-\gamma}.$
From Assumptions 3.3 and 3.4, we can easily see that $Q_{\max}\leq\frac{1}{1-\gamma}$.
4 Simultaneous double Q-learning (SDQ)
4.1 Algorithm
In this paper, we consider the following modified double $Q$-learning, called simultaneous double Q-learning (SDQ):
$Q_{k+1}^{A}(s_{k},a_{k})$ $=Q_{k}^{A}(s_{k},a_{k})+\alpha_{k}\{r_{k}+\gamma Q_{k}^{A}(s_{k+1},\text{argmax}_{a\in B}Q_{k}^{B}(s_{k},a))-Q_{k}^{A}(s_{k},a_{k})\},$
$Q_{k+1}^{B}(s_{k},a_{k})$ $=Q_{k}^{B}(s_{k},a_{k})+\alpha_{k}\{r_{k}+\gamma Q_{k}^{B}(s_{k+1},\text{argmax}_{a\in A}Q_{k}^{A}(s_{k},a))-Q_{k}^{B}(s_{k},a_{k})\}.$ (6)
The first difference between the original double $Q$-learning and SDQ is the role of each $Q$-estimator in the update. In the original double $Q$-learning, an optimal action is selected from the same $Q$-estimator, and it employs the other $Q$-estimator for bootstrapping. On the other hand, in the proposed version, an optimal action is selected from the other $Q$-estimator, and it employs the same $Q$-estimator for bootstrapping. This modification enables applying the switching system framework from [12], overcoming the difficulty caused by the switched order of $Q_{A}$ and $Q_{B}$ in original double $Q$-learning, while retaining its advantage of reducing overestimation bias.
The other difference is in the Bernoulli variable. Unlike the standard double $Q$-learning, which uses a Bernoulli variable for the $Q$-estimator selection, the modified version updates the two $Q$-estimators synchronously, which can potentially speed up the convergence. However, we note that our analysis can also include the Bernoulli random selection as in the original form without major changes in the finite-time error analysis. Besides, a potential issue that arises by eliminating the random $Q$-estimator selection is that if initially $Q^{A}_{0}=Q^{B}_{0}$, then $Q^{A}_{k}=Q^{B}_{k}$ for all $k\geq 0$. In this case, (4.1) is reduced to the standard $Q$-learning. To bypass the issue for implementation, one simple approach is to randomly initialize $Q^{A}_{0}$ and $Q^{B}_{0}$ so that $Q^{A}_{0}\neq Q^{B}_{0}$.
To demonstrate its effect, let us consider the example in Figure 1(left) (adopted from [1], Ch. 6.7). As observed in the result, SDQ initially suffers from overestimation of $Q$-values due to the random initialization of $Q$-estimators with uniform distribution. However, this bias is quickly mitigated. And it achieves similar result to the original double $Q$-learning as shown in Figure 1(right).
[…]
4.2 Experiment
In this section, we empirically compare SDQ with original double $Q$-learning on the environments, FrozenLake, CliffWalking, and Taxi in OpenAI Gym, and demonstrate that SDQ outperforms the standard double $Q$-learning. We utilize the epsilon-greedy exploration with the constant step-size $\alpha=0.01$ and discount factor $\gamma=0.99$ for both algorithms. $Q$-estimators for SDQ are initialized randomly with the uniform distribution, and intel core i7-11800h is used for all experiments. To account for the sparse rewards in FrozenLake, we evaluate the average episodic reward after applying a moving window of size $100$ for each episode. This approach smooths the reward signal, as the agent only receives a $+1$ reward upon successful completion. For all experiments, the results are averaged over $30$ independent runs. For comparative analysis, the baseline algorithms from https://github.com/ShangtongZhang/reinforcement-learning-an-introduction/tree/master/chapter06 are borrowed. As can be seen from Figure 2, SDQ exhibits better performances over the standard double $Q$-learning.
(a) FrozenLake-v0
(b) CliffWalking-v0
(c) Taxi-v3
Figure 2: Comparison of experiment results: Simultaneous Double $Q$-Learning (SDQ) vs. Double $Q$-Learning
4.3 Finite-time error bounds
In this subsection, we present finite-time error bounds for SDQ. Through the analysis given in this paper, we can derive a finite-time error bound given below.
Theorem 4.1.
For any $k\geq 0$, we have the following error bound:
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{120\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\rho^{k-4}k^{4}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}.$ (7)
The same bound holds for $Q_{k}^{B}-Q^{*}$.
The proof is given in Section C.5 of Appendix. The bound in (7) can be converted to more interpretable form presented below.
Corollary 4.2.
For any $k\geq 0$, we have the following error bound:
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{120\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\frac{\rho^{-4}(-8)^{4}}{(\ln(\rho))^{4}}\rho^{\frac{-4}{\ln(\rho)}}\rho^{k/2}.$ (8)
The same bound holds for $Q_{k}^{B}-Q^{*}$.
The proof is given in Section C.6 of Appendix. The remaining parts of the paper are devoted to brief sketches of the proofs.
5 Framework for convergence analysis of SDQ
5.1 Switching system model
In this subsection, we introduce a switching system model of SDQ in (4.1). First of all, using the notation introduced in Section 3.4, the modified update in (4.1) can be compactly written as
$Q_{k+1}^{A}=Q_{k}^{A}+\alpha_{k}(DR+\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-DQ_{k}^{A}+w_{k}^{A}),$
$Q_{k+1}^{B}=Q_{k}^{B}+\alpha_{k}(DR+\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}-DQ_{k}^{B}+w_{k}^{B}),$ (9)
where
$w_{k}^{A}$ $=(e_{a_{k}}\otimes e_{s_{k}})r_{k}+\gamma(e_{a}\otimes e_{s})(e_{s_{k}^{\prime}})^{T}\Pi_{Q_{k}^{B}}Q_{k}^{A}-(e_{a_{k}}\otimes e_{s_{k}})(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}$
$\quad-(DR+\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-DQ_{k}^{A}),$
$w_{k}^{B}$ $=(e_{a_{k}}\otimes e_{s_{k}})r_{k}+\gamma(e_{a}\otimes e_{s})(e_{s_{k}^{\prime}})^{T}\Pi_{Q_{k}^{A}}Q_{k}^{B}-(e_{a_{k}}\otimes e_{s_{k}})(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}$
$\quad-(DR+\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}-DQ_{k}^{B}).$ (10)
Next, using the optimal Bellman equation $(\gamma DP\Pi_{Q^{*}}-D)Q^{*}+DR=0$ with (5.1), one can obtain
$Q_{k+1}^{A}-Q^{*}$ $=(I-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{A}-\gamma P\Pi_{Q^{*}}Q^{*}\}+\alpha w_{k}^{A},$
$Q_{k+1}^{B}-Q^{*}$ $=(I-\alpha D)(Q_{k}^{B}-Q^{*})+\alpha D\{\gamma P\Pi_{Q_{k}^{A}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}\}+\alpha w_{k}^{B},$ (11)
which is a linear switching system with an extra affine terms, $\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{A}-\gamma P\Pi_{Q^{*}}Q^{*}\}$ and $\alpha D\{\gamma P\Pi_{Q_{k}^{A}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}\}$, and the stochastic noises, $w_{k}^{A}$ and $w_{k}^{B}$  [12].
The main difficulty in analysing the system is due to the extra affine term and the stochastic noise, without which the finite-time analysis is straightforward, as stability of the system matrix can be analyzed directly. However, with the affine term, the analysis becomes more challenging. To address this difficulty, we introduce the lower and upper comparison systems as in [12], which enable easier analysis.
5.2 Upper comparison system
Let us first consider the upper comparison system
$Q_{k+1}^{A_{U}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)(Q_{k}^{A_{U}}-Q^{*})+\alpha w_{k}^{A},\quad Q_{0}^{A_{U}}-Q^{*}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$
$Q_{k+1}^{B_{U}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q_{k}^{A}}-\alpha D)(Q_{k}^{B_{U}}-Q^{*})+\alpha w_{k}^{B},\quad Q_{0}^{B_{U}}-Q^{*}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert}.$ (12)
The above systems are switching systems, which have system matrices $I+\alpha\gamma DP\Pi_{Q_{k}^{A}}-\alpha D$ and $I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D$. These matrices switch according to the changes of $Q_{k}^{A}$ and $Q_{k}^{B}$.
We can prove that the trajectory of the upper comparison system bounds that of the original system from above.
Proposition 5.1.
Suppose that $Q_{0}^{A_{U}}-Q^{*}\geq Q_{0}^{A}-Q^{*}$ and $Q_{0}^{B_{U}}-Q^{*}\geq Q_{0}^{B}-Q^{*}$ hold, where $\geq$ is the element-wise inequality. Then, we have
$Q_{k}^{A_{U}}-Q^{*}\geq Q_{k}-Q^{*},\quad Q_{k}^{B_{U}}-Q^{*}\geq Q_{k}-Q^{*},$
for all $k\geq 0$.
The proof is in Section C.1 of Appendix.
5.3 Lower comparison system
Let us consider the lower comparison system
$Q_{k+1}^{A_{L}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$
$\quad+\alpha w_{k}^{A},\quad Q_{0}^{A_{L}}-Q^{*}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$
$Q_{k+1}^{B_{L}}-Q^{*}$ $=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{B_{L}}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})(Q_{k}^{A}-Q_{k}^{B})$
$\quad+\alpha w_{k}^{B},\quad Q_{0}^{B_{L}}-Q^{*}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$ (13)
where the stochastic noises $w_{k}^{A}$ and $w_{k}^{B}$ are identical to the original system (5.1).
As before, we can prove that the trajectory of the lower comparison system bounds that of the original system from below.
Proposition 5.2.
Suppose that $Q_{0}^{A_{L}}-Q^{*}\leq Q_{0}^{A}-Q^{*}$ and $Q_{0}^{B_{L}}-Q^{*}\leq Q_{0}^{B}-Q^{*}$ hold, where $\leq$ is the element-wise inequality. Then, we have
$Q_{k}^{A_{L}}-Q^{*}\leq Q_{k}^{A}-Q^{*},\quad Q_{k}^{B_{L}}-Q^{*}\leq Q_{k}^{B}-Q^{*},$
for all $k\geq 0$.
The proof is in Section 5.2 of Appendix. The lower comparison system (5.3) can be seen as a linear system with the states $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$ and the system matrix $I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D$. Moreover, it also includes the extra terms, $\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$ and $\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})(Q_{k}^{A}-Q_{k}^{B})$, which can be seen as external disturbances. To derive a finite-time error bound, one needs to establish bounds on the error $Q_{k}^{A}-Q_{k}^{B}$ first. Therefore, in the next subsections, we introduce an error system.
5.4 Error system
Let us consider the error system with the state $Q_{k}^{\text{err}}\coloneqq Q_{k}^{A}-Q_{k}^{B}$
$Q_{k+1}^{\text{err}}$ $=(I-\alpha D)Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}+\alpha w_{k}^{A}-\alpha w_{k}^{B},\quad Q_{0}^{\text{err}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$ (14)
which can be obtained by subtracting the switching system model of $Q_{k}^{B}$ from that of $Q_{k}^{A}$ in (5.1). The error system (14) can be seen as a linear system with the states $Q_{k}^{\text{err}}$ and the system matrix $I-\alpha D$. Moreover, it includes extra affine term $\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}$.
In the lower comparison system (5.3), the extra terms, $\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})Q_{k}^{\text{err}}$ and $\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})Q_{k}^{\text{err}}$, make it hard to analyze the finite-time error bounds of the lower comparison system compared to the original $Q$-learning [11, 12], where the lower comparison system is a linear system without the disturbance terms. To circumvent this difficulty, we will first prove that the error system $Q_{k}^{\text{err}}$ in the disturbance parts vanishes as $k\to\infty$. Intuitively, this implies that as the disturbance vanishes, and the lower comparison system converges to a pure stochastic linear system.
However, the error system in (14) has the affine term $\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}$ similar to the original double $Q$-learning or $Q$-learning. Therefore, one can imagine that its convergence can be proved using similar techniques as in the $Q$-learning analysis [11, 12, 13]. In particular, one can derive the upper and lower comparison systems of the error system, and then prove the overall convergence using the fact that the lower comparison system is linear. However, for the error system (14), similar ideas cannot be applied because the lower comparison system of the error system is a switching system. For this reason, we will use a different approach.
In particular, we will first consider an upper comparison system of the error system, and then, derive a lower comparison system of the upper comparison system, which is linear. Then, we will consider a lower comparison system of the error system. Because the lower comparison system is a switching system, we will derive a subtraction system which can be obtained by subtracting the error lower comparison system from the error upper comparison system.
6 Analysis process for convergence of SDQ
6.1 Overall plans
The overall flow of the proposed analysis is given in Figure 3. The texts highlighted with blue indicate the dynamic systems we will deal with for our analysis. The red arrows represent the directions we will follow for the proof. The overall process is summarized as follows: Step 1: The finite-time error bound of $Q_{k}^{\text{err}_{UL}}$ is obtained by using its corresponding linear system structure. Then, based on the error bound on $Q_{k}^{\text{err}_{\text{UL}}}$, the finite-time error bound on $Q_{k}^{\text{err}_{U}}$ can be derived. Step 2: following similar lines as in Step 1, one can derive the error bound on $Q_{k}^{\text{err}}$ based on the error bound on $Q_{k}^{\text{err}_{U}}$ and $Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}$. Step 3: Using the error bound on $Q_{k}^{\text{err}}$ and the linear structures of $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$, the finite-time error bounds on $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$ can be derived. Step 4: By obtaining a subtraction system which can be obtained by subtracting the error lower comparison system from the error upper comparison system, the convergence of $Q_{k}^{A_{U}}-Q^{*}$ and $Q_{k}^{B_{U}}-Q^{*}$ can be shown. Step 5: Using the previous results, we can obtain a finite-time error bound on the iterates of SDQ. These steps will be detailed in Appendix.
Figure 3: Overall flow of the proposed analysis
7 Discussion and Conclusion
In this paper, we present a novel variant of double $Q$-learning, called simultaneous double $Q$-learning (SDQ), which empirically achieves faster convergence than the original double $Q$-learning by eliminating the random selection of double $Q$-estimators. Moreover, SDQ introduces a novel switching system perspective by swapping the roles of the two $Q$-estimators. By leveraging switching system representations, we establish novel finite-time expected error bounds for SDQ. It is also important to note that while the random $Q$-estimator selection step is eliminated in SDQ, it needs for the additional initialization step for the $Q$-estimator, which can affect the overall convergence properties. Therefore, future researches can aim to address this initialization effect.
Supplementary material
Appendix A Convergence of stochastic linear system
To aid in understanding the intricacies of the convergence of simultaneous double $Q$-learning(SDQ), as explained in Sections B and C, let us consider the following stochastic linear system, which offers a helpful conceptual framework:
$x_{k+1}=Ax_{k}+\alpha v_{k},\quad x_{0}\in\mathbb{R}^{n},\quad k\in\{0,1,...\},$ (15)
where $x_{k}\in\mathbb{R}^{n}$ is the state, and $A$ is system matrix, and $\alpha v_{k}$ is the stochastic noise with a constant $\alpha\in(0,1)$. Here, we will first investigate a finite-time error analysis of the state of (15), and it will be used in the proof of SDQ. To this end, let us assume that the covariance matrix of $v_{k}$ is bounded as $\mathbb{E}[v_{k}v_{k}^{T}]\leq V_{\text{max}}$ and $V_{\text{max}}>0$. Then, it can be proved that the maximum eigenvalue of $\mathbb{E}[v_{k}v_{k}^{T}]$ can be bounded by $V_{\text{max}}$.
Lemma 1.
[13] The maximum eigenvalue of $\mathbb{E}[v_{k}v_{k}^{T}]$ is bounded as
$\lambda_{\text{max}}(\mathbb{E}[v_{k}v_{k}^{T}])\leq V_{\text{max}}$
for all $k\geq 0$, where $V_{\text{max}}>0$ is from our assumption.
Proof.
The proof is completed by noting $\lambda_{\text{max}}(\mathbb{E}[v_{k}v_{k}^{T}])\leq\text{tr}(\mathbb{E}[v_{k}v_{k}^{T}])=\mathbb{E}[\text{tr}(v_{k}v_{k}^{T})]=\mathbb{E}[v_{k}^{T}v_{k}]\leq V_{\text{max}},$ where the last inequality comes from our assumption and the second equality uses the fact that the trace is a linear function. This completes the proof. ∎
Moreover, let us assume that the system matrix $A$ is also bounded.
Assumption A.1.
The system matrix $A$ satisfies $A\leq\rho$ with some constant $\rho\in(0,1)$.
As a next step, we investigate how the auto-correlation matrix $\mathbb{E}[x_{k}x_{k}^{T}]$ propagates over the time. Thus, one can consider the auto-correlation matrix of the state recursively calculated as follows:
$\mathbb{E}[x_{k+1}x_{k+1}^{T}]=A\mathbb{E}[x_{k}x_{k}^{T}]A^{T}+\alpha^{2}V_{k},$
where $\mathbb{E}[v_{k}v_{k}^{T}]=V_{k}$. Defining $X_{k}\coloneqq\mathbb{E}[x_{k}x_{k}^{T}],k\geq 0$, the above recursion can be written by
$X_{k+1}=AX_{k}A^{T}+\alpha^{2}V_{k},\quad\forall k\geq 0.$
To prove the convergence of (15), we first establish a bound on the trace of $X_{k}$.
Lemma 2.
[13] We have the following bound:
$\text{tr}(X_{k})\leq\frac{9n^{2}\alpha}{d_{\text{min}}(1-\gamma)^{3}}+\lVert x_{0}\rVert^{2}_{2}n^{2}\rho^{2k}$
Proof.
We first bound $\lambda_{\text{max}}(X_{k})$ as follows:
$\begin{split}\lambda_{\text{max}}(X_{k})&\leq\alpha^{2}\sum_{i=0}^{k-1}\lambda_{\text{max}}(A^{i}V_{k-i-1}(A^{T})^{i})+\lambda_{\text{max}}(A^{k}X_{0}(A^{T})^{k})\\ &\leq\alpha^{2}\sup_{j\geq 0}\lambda_{\text{max}}\sum_{i=0}^{k-1}\lambda_{\text{max}}(A^{i}(A^{T})^{i})+\lambda_{\text{max}}(X_{0})\lambda_{\text{max}}(A^{k}(A^{T})^{k})\\ &=\alpha^{2}\sup_{j\geq 0}\lambda_{\text{max}}(V_{j})\sum_{i=0}^{k-1}\lVert A^{i}\rVert_{2}^{2}+\lambda_{\text{max}}(X_{0})\lVert A^{k}\rVert_{2}^{2}\\ &\leq\alpha^{2}V_{\text{max}}n\sum_{i=0}^{k-1}\lVert A^{i}\rVert_{\infty}^{2}+n\lambda_{\text{max}}(X_{0})\lVert A^{k}\rVert_{\infty}^{2}\\ &\leq\alpha^{2}V_{\text{max}}n\sum_{i=0}^{k-1}\rho^{2i}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\\ &\leq\alpha^{2}V_{\text{max}}n\lim_{k\rightarrow\infty}\sum_{i=0}^{k-1}\rho^{2i}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\\ &\leq\frac{\alpha^{2}V_{\text{max}}n}{1-\rho^{2}}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\\ &\leq\frac{\alpha^{2}V_{\text{max}}n}{1-\rho}+n\lambda_{\text{max}}(X_{0})\rho^{2k}\end{split}$
where the first inequality is due to $A^{i}V_{k-i-1}(A^{T})^{i}\succeq 0$ and $A^{k}X_{0}(A^{T})^{k}\succeq 0$, the third inequality comes from Lemma 1, $\lVert\cdot\rVert_{2}\leq\sqrt{n}\lVert\cdot\rVert_{\infty}$, the fourth inequality is due to A.1, and the sixth and last inequalities come from $\rho\in(0,1)$. On the other hand, since $X_{k}\succeq 0$, the diagonal elements are nonnegative. Therefore, we have $\text{tr}(X_{k})\leq n\lambda_{\text{max}}(X_{k})$. Combining the last two inequalities leads to
$\text{tr}(X_{k})\leq n\lambda_{\text{max}}(X_{k})\leq\frac{\alpha^{2}V_{\text{max}}n^{2}}{1-\rho}+n^{2}\lambda_{\text{max}}(X_{0})\rho^{2k}$
Moreover, noting the inequality $\lambda_{\text{max}}(X_{0})\leq\text{tr}(X_{0})=\text{tr}(x_{0}x_{0}^{T})=\lVert x_{0}\rVert_{2}^{2}$, and plugging $\rho=1-\alpha d_{\text{min}}(1-\gamma)$ into $\rho$ in the last inequality, one gets the desired conclusion. ∎
Now, we are ready to present a finite-time bound on the state $x_{k}$ of (15).
Theorem A.2.
[13] For any $k\geq 0$, we have
$\mathbb{E}\left[\lVert x_{k}\rVert_{2}\right]$ $\leq\frac{3\alpha^{1/2}n}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+n\lVert x_{0}\rVert_{2}\rho^{k}$ (16)
Proof.
Noting the relations
$\begin{split}\mathbb{E}[\lVert x_{k}\rVert_{2}^{2}]=\mathbb{E}[\lVert x_{k}^{T}x_{k}\rVert]=\mathbb{E}[\lVert\text{tr}(x_{k}^{T}x_{k})\rVert]=\mathbb{E}[\lVert\text{tr}(x_{k}x_{k}^{T})\rVert]=\mathbb{E}[\text{tr}(X_{k})],\end{split}$
and using the bound in Lemma 2, one gets
$\mathbb{E}\left[\lVert x_{k}\rVert_{2}^{2}\right]$ $\leq\frac{9\alpha n^{2}}{d_{\text{min}}(1-\gamma)^{3}}+n^{2}\lVert x_{0}\rVert_{2}^{2}\rho^{2k}$
Taking the square root on both side of the last inequality, using the subadditivity of the square root function, the Jensen inequality, and the concavity of the square root function, we have the desired conclusion. ∎
Note that the result in Theorem A.2 will be used in our main analysis of SDQ. In particular, we will use the following form of the state of the system:
$x_{i}=A^{i}x_{0}+\sum_{j=0}^{i-1}\alpha A^{(i-1)-j}v_{j},$ (17)
which can be obtained by summing the recursion in (15) from $k=0$ to $k=i$. Based on the above expression, Theorem A.2 can be presented different form as follows.
Corollary A.3.
For any $k\geq 0$, we have
$\mathbb{E}\biggl{[}\biggl{\lVert}A^{k}x_{0}+\sum_{j=0}^{k-1}\alpha A^{(k-1)-j}v_{j}\biggr{\rVert}_{2}\biggl{]}\leq\frac{3\alpha^{1/2}n}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+n\lVert x_{0}\rVert_{2}A^{k}$
Proof.
The proof can be done directly from (17) and Theorem A.2. ∎
Appendix B Detailed analysis result of the convergence of SDQ ($Q_{k}^{\text{err}}$ part)
To establish a finite-time error bound in this paper, the main challenge is to establish a bound of the error system $Q_{k}^{\text{err}}$. The overall analysis strategy is presented in Section 6.1 briefly. We derive the convergence of $Q_{k}^{\text{err}}$ using the following two steps:
  • Step 1: A finite-time error bound of $Q_{k}^{\text{err}_{UL}}$ is obtained by using its corresponding linear system structure. Then, based on the error bound on $Q_{k}^{\text{err}_{UL}}$, a finite-time error bound on $Q_{k}^{\text{err}_{U}}$ can be derived.
  • Step 2: Next, following similar lines as in Step 1, one can derive an error bound on $Q_{k}^{\text{err}}$ based on the error bound on $Q_{k}^{\text{err}_{U}}$ and $Q_{k}^{\text{err}_{L}}$.
In this section, we present a detailed analysis process. To establish the groundwork for our proof, we first introduce an auxiliary Lemma demonstrating the boundedness of the system matrix.
Lemma 1.
[13] For any $Q\in{\mathbb{R}}^{\lvert S\rvert\lvert A\rvert}$, we have
$\lVert A_{Q}\rVert_{\infty}\leq\rho,$
where the matrix norm $\lVert A\rVert_{\infty}\coloneqq max_{1\leq i\leq m}\sum_{j=1}^{n}\lvert A_{ij}\rvert$ and $A_{ij}$ is the element of $A$ in $i$-th row and $j$-th column.
Proof.
Note the following identities
$\begin{split}\sum_{j}\lvert[A_{Q}]_{ij}\rvert&=\sum_{j}\lvert[I-\alpha D+\alpha\gamma DP\Pi_{Q}]_{ij}\rvert\\ &=[I-\alpha D]_{ii}+\sum_{j}[\alpha\gamma DP\Pi_{Q}]_{ij}\\ &=1-\alpha[D]_{ii}+\alpha\gamma[D]_{ii}\sum_{j}[P\Pi_{Q}]_{ij}\\ &=1-\alpha[D]_{ii}+\alpha\gamma[D]_{ii}\\ &=1+\alpha[D]_{ii}(\gamma-1),\end{split}$
where the second line is due to the fact that $A_{Q}$ is a non-negative matrix. Taking the maximum over $i$, we have
$\begin{split}\lVert A_{Q}\rVert_{\infty}&=\max_{i\in\{1,2,...,\lvert S\rvert\lvert A\rvert\}}{1+\alpha[D]_{ii}(\gamma-1)}\\ &=1-\alpha\min_{(s,a)\in\mathit{S}\times\mathit{A}}d(s,a)(1-\gamma)\\ &=\rho,\end{split}$
which completes the proof. ∎
As the first step, we present a convergence analysis of $Q_{k}^{\text{err}_{U}}$ in next subsection.
B.1 Convergence of $Q_{k}^{\text{err}_{U}}$
Let us write the error upper comparison system $Q_{k}^{\text{err}_{U}}$ as follows:
$Q_{k+1}^{\text{err}_{U}}=(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B},\quad Q_{0}^{\text{err}_{U}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$ (18)
where the stochastic noises, $w_{k}^{A}$ and $w_{k}^{B}$, are identical to those of the original system in (5.1). In the following proposition, we prove that $Q_{k}^{\text{err}_{U}}$ upper bounds $Q_{k}^{\text{err}}$.
Proposition B.1.
Suppose $Q_{0}^{\text{err}_{U}}\geq Q_{0}^{\text{err}}$, where $\geq$” is used as the element-wise inequality. Then, we have
$Q_{k}^{\text{err}_{U}}\geq Q_{k}^{\text{err}}$
for all $k\geq 0.$
Proof.
The proof is completed by an induction argument. Suppose that $Q_{i}^{\text{err}_{U}}\geq Q_{i}^{\text{err}}$ holds for $0\leq i\leq k$. Then, it follows that
$\begin{split}Q_{k+1}^{\text{err}}&=Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}-\alpha DQ_{k}^{\text{err}}\\ &\quad+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}-\alpha DQ_{k}^{\text{err}}\\ &\quad+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{\text{err}}-\alpha DQ_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=(I+\alpha\gamma DP\Pi_{Q_{k}^{A}}-\alpha D)Q_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}}}-\alpha D)Q_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=Q_{k+1}^{\text{err}_{U}},\end{split}$
where the first inequality is due to $\Pi_{Q_{k}^{A}}Q_{k}^{A}\geq\Pi_{Q_{k}^{B}}Q_{k}^{A}$ and the second inequality is due to $\Pi_{Q_{k}^{A}-Q_{k}^{B}}Q_{k}^{\text{err}}\geq\Pi_{Q_{k}^{A}}Q_{k}^{\text{err}}$, respectively, and the third inequality is due to the hypothesis $Q_{k}^{\text{err}_{U}}\geq Q_{k}^{\text{err}}$ and the fact that the matrix $I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}}}-\alpha D$ is nonnegative, i.e., all elements are nonnegative by Lemma 1. Therefore, $Q_{k+1}^{\text{err}_{U}}\geq Q_{k+1}^{\text{err}}$ holds, and the proof is completed by induction. ∎
To prove the convergence of $Q_{k}^{\text{err}_{U}}$, we consider another comparison system $Q_{k}^{\text{err}_{UL}}$ which is a lower comparison system of $Q_{k}^{\text{err}_{U}}$. In the following subsection, a convergence analysis of $Q_{k}^{\text{err}_{UL}}$ is presented.
B.2 Convergence of $Q_{k}^{\text{err}_{UL}}$
Let us write the $Q_{k}^{\text{err}_{UL}}$, which has the following form:
$Q_{k+1}^{\text{err}_{UL}}=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)Q_{k}^{\text{err}_{UL}}+\alpha w_{k}^{A}-\alpha w_{k}^{B},$ (19)
where the stochastic noises $w_{k}^{A}$ and $w_{k}^{B}$ are identical to those in the original system (5.1). Note that the system is the lower comparison system of the upper comparison system corresponding to $Q_{k}^{\text{err}}$.
In the following proposition, we prove that $Q_{k}^{\text{err}_{UL}}$ is a lower comparison system of $Q_{k}^{\text{err}_{U}}$.
Proposition B.2.
Suppose $Q_{0}^{\text{err}_{U}}\geq Q_{0}^{\text{err}_{UL}}$, where $\geq$” is used as the element-wise inequality. Then, we have
$Q_{k}^{\text{err}_{U}}\geq Q_{k}^{\text{err}_{UL}}$
for all $k\geq 0$.
Proof.
The proof is completed by an induction argument. Suppose that $Q_{i}^{\text{err}_{U}}\geq Q_{i}^{\text{err}_{UL}}$ holds for $0\leq i\leq k$. Then, it follows from (18) that
$\begin{split}Q_{k+1}^{\text{err}_{U}}&=(I+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\geq(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)Q_{k}^{\text{err}_{U}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\geq(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)Q_{k}^{\text{err}_{UL}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=Q_{k+1}^{\text{err}_{UL}},\end{split}$
where the first inequality is due to $\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{U}}\geq\Pi_{Q^{*}}Q_{k}^{\text{err}_{U}}$ and the second inequality is due to the hypothesis $Q_{k}^{\text{err}_{U}}\geq Q_{k}^{\text{err}_{UL}}$ and the fact that the matrix $I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D$ is nonnegative, i.e., all elements are nonnegative by Lemma 1. Therefore, $Q_{k+1}^{\text{err}_{U}}\geq Q_{k+1}^{\text{err}_{UL}}$ holds, and the proof is completed by induction.
The system (B.3) is a stochastic linear system with system matrix $I+\alpha\gamma DP\Pi_{Q_{k}^{*}}-\alpha D$ and noise $w_{k}^{A}-w_{k}^{B}$. To establish the convergence bound of this system, the same analysis approach as in Section A can be applied.
In particular, let us define $x_{k}\coloneqq Q_{k}^{\text{err}_{UL}}$ and $A\coloneqq I+\alpha\gamma DP\Pi_{Q_{k}^{*}}-\alpha D$. Then, the system (B.3) can be presented as the following stochastic linear system:
$x_{k+1}=Ax_{k}+\alpha(w_{k}^{A}-w_{k}^{B}),\quad x_{0}\in\mathbb{R}^{n},\quad\forall k\geq 0,$ (20)
where the noise term $w_{k}^{A}-w_{k}^{B}$ can be written as
$w_{k}^{A}-w_{k}^{B}=(e_{a_{k}}\otimes e_{s_{k}})(\delta_{k}^{A}-\delta_{k}^{B})-\gamma DP(\Pi_{Q^{B}_{k}}Q^{A}_{k}-\Pi_{Q^{A}_{k}}Q^{B}_{k})+D(Q_{k}^{A}-Q_{k}^{B}),\$
where
$\begin{split}w_{k}^{A}&=(e_{a_{k}}\otimes e_{s_{k}})r_{k}^{A}+\gamma(e_{a_{k}}\otimes e_{s_{k}})(e_{s_{k}^{\prime}})^{T}\Pi_{Q^{B}_{k}}Q^{A}_{k}-(e_{a_{k}}\otimes e_{s_{k}})(e_{a_{k}}\otimes e_{s_{k}})^{T}Q^{A}_{k}\\ &\quad-(DR+\gamma DP\Pi_{Q^{B}_{k}}Q^{A}_{k}-DQ_{k}^{A})\\ w_{k}^{B}&=(e_{a_{k}}\otimes e_{s_{k}})r_{k}^{B}+\gamma(e_{a_{k}}\otimes e_{s_{k}})(e_{s_{k}^{\prime}})^{T}\Pi_{Q^{A}_{k}}Q^{B}_{k}-(e_{a_{k}}\otimes e_{s_{k}})(e_{a_{k}}\otimes e_{s_{k}})^{T}Q^{B}_{k}\\ &\quad-(DR+\gamma DP\Pi_{Q^{A}}Q^{B}-DQ_{k}^{B})\\ \delta_{k}^{A}&\coloneqq r_{k}^{A}+\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{B}}Q_{k}^{A}-(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}\\ \delta_{k}^{B}&\coloneqq r_{k}^{B}+\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{A}}Q_{k}^{B}-(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}\end{split}$
To prove the convergence of $Q_{k}^{\text{err}_{UL}}$, we prove the boundedness of the noise term in (20).
The boundedness of $w_{k}^{A}-w_{k}^{B}$ in (20) is formally proved in the following lemma.
Lemma 2.
The noise term $w_{k}^{A}-w_{k}^{B}$ in (20) satisfies
$\mathbb{E}\left[(w_{k}^{A}-w_{k}^{B})^{T}(w_{k}^{A}-w_{k}^{B})\right]\leq\frac{16}{(1-\gamma)^{2}}\coloneqq W_{\text{max}}$
for all $k\geq 0$
Proof.
One can get the following bound on $\mathbb{E}\left[(w_{k}^{A}-w_{k}^{B})^{T}(w_{k}^{A}-w_{k}^{B})\right]$:
$\begin{split}&\mathbb{E}\left[(w_{k}^{A}-w_{k}^{B})^{T}(w_{k}^{A}-w_{k}^{B})\right]\\ =&\mathbb{E}\left[\lVert((e_{a_{k}}\otimes e_{s_{k}})(\delta_{k}^{A}-\delta_{k}^{B}))-(\gamma DP(\Pi_{Q^{B}_{k}}Q^{A}_{k}-\Pi_{Q^{A}_{k}}Q^{B}_{k})-D(Q_{k}^{A}-Q_{k}^{B}))\rVert_{2}^{2}\right]\\ =&\mathbb{E}\left[(\delta_{k}^{A}-\delta_{k}^{B})^{2}\right]-\lVert\gamma DP(\Pi_{Q^{B}_{k}}Q^{A}_{k}-\Pi_{Q^{A}_{k}}Q^{B}_{k})-D(Q_{k}^{A}-Q_{k}^{B}))\rVert_{2}^{2}\\ \leq&\mathbb{E}\left[(\delta_{k}^{A}-\delta_{k}^{B})^{2}\right]\\ =&\mathbb{E}\left[(r_{k}^{A}+\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{B}}Q_{k}^{A}-(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}-(r_{k}^{B}+\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{A}}Q_{k}^{B}-(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}))^{2}\right]\\ \leq&\mathbb{E}\left[(|r_{k}^{A}|+|\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{B}}Q_{k}^{A}|+|(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{A}|+(|r_{k}^{B}|+|\gamma(e_{s_{k}^{\prime}}^{T})\Pi_{Q_{k}^{A}}Q_{k}^{B}|+|(e_{a_{k}}\otimes e_{s_{k}})^{T}Q_{k}^{B}|))^{2}\right]\\ =&\frac{16}{(1-\gamma)^{2}}=W_{\text{max}}\end{split}$
Now, because (B.3) is a stochastic linear system, the analysis of a simple stochastic linear system from Section A can be applied directly. Then, we can get the upper bound of $Q_{k}^{\text{err}_{UL}}$ in the following lemma.
Lemma 3.
For any $k\geq 0$, we have
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{2}]\leq\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\rho^{k}$
Proof.
Noting the relations
$\begin{split}\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{2}^{2}]&=\mathbb{E}[\lVert(Q_{k}^{\text{err}_{UL}})^{T}(Q_{k}^{\text{err}_{UL}})\rVert]\\ &=\mathbb{E}[\lVert\text{tr}(Q_{k}^{\text{err}_{UL}})^{T}(Q_{k}^{err_{UL}})\rVert]\\ &=\mathbb{E}[\lVert\text{tr}((Q_{k}^{\text{err}_{UL}})(Q_{k}^{err_{UL}})^{T}\rVert]\\ &=\mathbb{E}[\lVert\text{tr}(X_{k})\rVert]\end{split}$
and using the bound in Lemma 2 and Lemma 2 lead to
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{2}^{2}]\leq\frac{16\alpha\lvert S\times A\rvert^{2}}{d_{\text{min}}(1-\gamma)^{3}}+\lvert S\times A\rvert^{2}\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}^{2}\rho^{2k}$
Taking the square root on both side of the last inequality, using the subadditivity of the square root function, the Jensen inequality, and the concavity of the square root function, we have the desired conclusion. ∎
To get the upper bound of $Q_{k+1}^{\text{err}_{U}}$, subtracting the lower comparison system $Q_{k+1}^{\text{err}_{UL}}$ from upper comparison system $Q_{k+1}^{\text{err}_{U}}$, the following form can be obtained:
$Q_{k+1}^{\text{err}_{U}}-Q_{k+1}^{\text{err}_{UL}}$ $=(I-\alpha D)(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}})+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{*}}Q_{k}^{\text{err}_{UL}})$
$=(I-\alpha D)(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}})$
$\quad+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{UL}}+\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{UL}}-\Pi_{Q_{k}^{*}}Q_{k}^{\text{err}_{UL}})$
$=(I-\alpha D+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}})(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}})$
$\quad+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}-\Pi_{Q_{k}^{*}})Q_{k}^{\text{err}_{UL}}$ (21)
Taking the $\infty$-norm and expectation on (B.2) yields the bound
$\mathbb{E}[\lVert Q_{k+1}^{\text{err}_{U}}-Q_{k+1}^{\text{err}_{UL}}\rVert_{\infty}]$ $\leq\lVert I-\alpha D+\alpha\gamma DP\Pi_{Q_{k}^{\text{err}_{U}}}\rVert_{\infty}\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$
$\quad+\lVert\alpha\gamma DP\rVert_{\infty}\lVert\Pi_{Q_{k}^{\text{err}_{U}}}-\Pi_{Q_{k}^{*}}\rVert_{\infty}\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$
$\leq\rho\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]+\alpha\gamma d_{\text{max}}\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$
$\leq\rho\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$
$\quad+\alpha\gamma d_{\text{max}}\biggl{(}\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\rho^{k}\biggl{)},$ (22)
where the second inequality is due to Lemma 1 and the last inequality is due to Lemma 3. Letting $Q_{0}^{\text{err}_{U}}=Q_{0}^{\text{err}_{UL}}$ in (B.3) and applying the inequality successively result in
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$ $\leq\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$ (23)
Using this result, we can obtain the bound of $\mathbb{E}\bigl{[}\lVert Q_{k}^{\text{err}_{U}}\rVert_{\infty}\bigr{]}$.
Thus, $Q_{k}^{\text{err}_{U}}$ satisfies
$\mathbb{E}\bigl{[}\lVert Q_{k}^{\text{err}_{U}}\rVert_{\infty}\bigr{]}$ $=\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}+Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$
$\leq\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]+\mathbb{E}[\lVert Q_{k}^{\text{err}_{UL}}\rVert_{\infty}]$
$\leq\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\rho^{k}$
$\leq\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k},$ (24)
where the second equality is due to (23) and Lemma 3 and the last inequality is due to the following fact $\lVert Q_{0}^{\text{err}_{UL}}\rVert_{2}\leq\lvert S\times A\rvert^{1/2}\lVert Q_{0}^{\text{err}_{UL}}\rVert_{\infty}\leq\lvert S\times A\rvert^{1/2}\frac{2}{1-\gamma}$. Because the upper comparison system bounds all trajectory that of original system, we use this bound as the upper bound of the original system.
B.3 CONVERGENCE OF $Q_{k}^{\text{err}_{L}}$
As the next step for the convergence analysis of $Q_{k}^{\text{err}}$, let us write the error lower comparison system $Q_{k}^{\text{err}_{L}}$ as follows:
$Q_{k+1}^{\text{err}_{L}}=(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)Q_{k}^{\text{err}_{L}}+\alpha w_{k}^{A}-\alpha w_{k}^{B},\quad Q_{0}^{\text{err}_{L}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$
where the stochastic noises, $w_{k}^{A}$ and $w_{k}^{B}$, are identical to those of the original system in (5.1). In the following proposition, we prove that $Q_{k}^{\text{err}_{L}}$ lower bounds $Q_{k}^{\text{err}}$.
Proposition B.3.
Suppose $Q_{0}^{\text{err}_{L}}\leq Q_{0}^{\text{err}}$, where $\leq$” is used as the element-wise inequality. Then, we have
$Q_{k}^{\text{err}_{L}}\leq Q_{k}^{\text{err}}$
for all $k\geq 0.$
Proof.
The proof is completed by an induction argument. Suppose that $Q_{i}^{\text{err}_{L}}\leq Q_{i}^{\text{err}}$ holds for $0\leq i\leq k$. Then, it follows that
$\begin{split}Q_{k+1}^{\text{err}}&=Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{A}}Q_{k}^{B}-\alpha DQ_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\geq Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{A}-\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{B}-\alpha DQ_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=Q_{k}^{\text{err}}+\alpha\gamma DP\Pi_{Q_{k}^{B}}Q_{k}^{\text{err}}-\alpha DQ_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)Q_{k}^{\text{err}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &\geq(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)Q_{k}^{\text{err}_{L}}+\alpha w_{k}^{A}-\alpha w_{k}^{B}\\ &=Q_{k+1}^{\text{err}_{L}},\end{split}$
where the first inequality is due to $\Pi_{Q_{k}^{B}}Q_{k}^{B}\geq\Pi_{Q_{k}^{B}}Q_{k}^{A}$, and the second inequality is due to the hypothesis $Q_{k}^{\text{err}_{L}}\leq Q_{k}^{\text{err}}$ and the fact that the matrix $I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D$ is nonnegative, i.e., all elements are nonnegative by Lemma 1. Therefore, $Q_{k+1}^{\text{err}_{L}}\leq Q_{k+1}^{\text{err}}$ holds, and the proof is completed by induction. ∎
The error lower comparison system switches according to the change of $Q_{k}^{B}$. So it is hard to analyze the stability of the lower comparison system in contrast to (B.3) which is linear system. To circumvent such a difficulty, we instead study an subtraction system by subtracting the error lower comparison system from the error upper comparison system as follows
$Q_{k+1}^{\text{err}_{U}}-Q_{k+1}^{\text{err}_{L}}$ $=(I-\alpha D)(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}})+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{B}}Q_{k}^{\text{err}_{L}})$
$=(I-\alpha D)(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}})$
$\quad+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{{B}}}Q_{k}^{\text{err}_{L}}+\Pi_{Q_{k}^{{B}}}Q_{k}^{\text{err}_{U}}-\Pi_{Q_{k}^{B}}Q_{k}^{\text{err}_{U}})$
$=(I-\alpha D+\alpha\gamma DP\Pi_{Q_{k}^{B}})(Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}})+\alpha\gamma DP(\Pi_{Q_{k}^{\text{err}_{U}}}-\Pi_{Q_{k}^{B}})Q_{k}^{\text{err}_{U}},$ (25)
Here the stochastic noise is canceled out in the error system. The key insight is as follows: if we can prove the stability of the subtraction system, i.e., $Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}\rightarrow 0$ as $k\rightarrow\infty$, then since $Q_{k}^{\text{err}_{U}}\rightarrow 0$ we have $Q_{k}^{\text{err}_{L}}\rightarrow 0$
Taking the $\infty$-norm and expectation on (B.3) yields the bound
$\mathbb{E}[\lVert Q_{k+1}^{\text{err}_{U}}-Q_{k+1}^{\text{err}_{L}}\rVert_{\infty}]$ $\leq\lVert I-\alpha D+\alpha\gamma DP\Pi_{Q_{k}^{{B}}}\rVert_{\infty}\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}\rVert_{\infty}]$
$\quad+\lVert\alpha\gamma DP\rVert_{\infty}\lVert\Pi_{Q_{k}^{\text{err}_{U}}}-\Pi_{Q_{k}^{B}}\rVert_{\infty}\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}\rVert_{\infty}]$
$\leq\rho\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}\rVert_{\infty}]+\alpha\gamma d_{\text{max}}\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}\rVert_{\infty}]$
$\leq\rho\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}\rVert_{\infty}]$
$\quad+\alpha\gamma d_{\text{max}}\biggl{(}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}\biggl{)},$ (26)
where the second inequality is due to Lemma 1 and the last inequality is due to Lemma 3. Letting $Q_{0}^{\text{err}_{U}}=Q_{0}^{\text{err}_{L}}$ in (B.3) and applying the inequality successively result in
$\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}-Q_{k}^{\text{err}_{L}}\rVert_{\infty}]$ $\leq\rho^{k}\mathbb{E}[\lVert Q_{0}^{\text{err}_{U}}-Q_{0}^{\text{err}_{L}}\rVert_{\infty}]+\alpha\gamma d_{\text{max}}\biggl{(}\frac{\rho^{k}}{1-\rho}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+\frac{\rho^{k}}{1-\rho}\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\rho^{k-2}\frac{(k-1)(k-2)}{2}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+k\rho^{k-1}\frac{2\lvert S\times A\rvert^{3/2}}{1-\gamma}\biggl{)}$
$=\rho^{k}\frac{4\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\rho^{k}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+\rho^{k-2}(k-1)(k-2)\frac{\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+k\rho^{k-1}\frac{2\lvert S\times A\rvert^{3/2}\alpha\gamma d_{\text{max}}}{1-\gamma},$ (27)
where the equality is due to Definition 3.5.
B.4 CONVERGENCE OF $Q_{k}^{\text{err}}$
By using upper comparison system and upper-lower comparison system and lower comparison system corresponding to the error system, one can derive the finite-time bound of $Q_{k}^{\text{err}}$.
Lemma 4.
For any $k\geq 0$, we have
$\mathbb{E}\left[\lVert Q_{k}^{\text{err}}\rVert_{\infty}\right]$ $\leq\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\frac{8\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{4\rho^{k-2}k^{2}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\frac{4k\rho^{k-1}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}.$ (28)
Proof.
We can get the bound of $Q_{k}^{\text{err}}$ as follows
$\mathbb{E}\left[\lVert Q_{k}^{\text{err}}\rVert_{\infty}\right]$ $=\mathbb{E}[\lVert Q_{k}^{\text{err}}-Q_{k}^{\text{err}_{U}}+Q_{k}^{\text{err}_{U}}\rVert_{\infty}]$
$\leq\mathbb{E}[\lVert Q_{k}^{\text{err}}-Q_{k}^{\text{err}_{U}}\rVert_{\infty}]+\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}\rVert_{\infty}]$
$\leq\mathbb{E}[\lVert Q_{k}^{\text{err}_{L}}-Q^{\text{err}_{U}}\rVert_{\infty}]+\mathbb{E}[\lVert Q_{k}^{\text{err}_{U}}\rVert_{\infty}]$
$\leq\rho^{k}\frac{4\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\rho^{k}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+\rho^{k-2}(k-1)(k-2)\frac{\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}+\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\quad+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}$ (29)
Then we group some terms of (B.4) as
$\rho^{k}\frac{4\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\rho^{k}\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}\leq\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}$
$\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}\leq\frac{8\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}},$
$\rho^{k-2}(k-1)(k-2)\frac{\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}$
$\leq\frac{4\rho^{k-2}k^{2}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)},$
$\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}+k\rho^{k-1}2\alpha\gamma d_{\text{max}}\frac{\lvert S\times A\rvert^{3/2}}{1-\gamma}\leq\frac{4k\rho^{k-1}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
Then we can get a simplified form as
$\mathbb{E}\left[\lVert Q_{k}^{\text{err}}\rVert_{\infty}\right]$ $\leq\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}+\frac{8\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{4\rho^{k-2}k^{2}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\frac{4k\rho^{k-1}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
Appendix C Detailed analysis result of convergence of SDQ (Remaining part)
In this section, convergence of SDQ will be studied based on the results in Section B. In Section B, a bound on $Q_{k}^{\text{err}}$ has been obtained in Lemma 4. To compete the proof, the following three steps remain:
  • Step 3: Using the bound on $Q_{k}^{\text{err}}$ and the linear structures of $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$, a finite-time error bounds on $Q_{k}^{A_{L}}-Q^{*}$ and $Q_{k}^{B_{L}}-Q^{*}$ can be derived.
  • Step 4: By obtaining a subtraction system which can be obtained by subtracting the lower comparison system from upper comparison system, the convergence of $Q_{k}^{A_{U}}-Q^{*}$ and $Q_{k}^{B_{U}}-Q^{*}$ can be shown.
  • Step 5: Next, combining the result from Step 4 with the upper comparison system $Q_{k}^{A_{U}}-Q^{*}$ and $Q_{k}^{B_{U}}-Q^{*}$, we can finally obtain the finite-time error bound on the iterates of SDQ.
C.1 Proof of Proposition 5.1 (Upper comparison system)
Using the dynamic system equation (5.1), we have
$\begin{split}Q_{k+1}^{A}-Q^{*}&=Q_{k}^{A}-Q^{*}+\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{A}-\gamma P\Pi_{Q^{*}}Q^{*}-Q_{k}^{A}+Q^{*}\}+\alpha w_{k}^{A}\\ &\leq Q_{k}^{A}-Q^{*}+\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{A}-\gamma P\Pi_{Q_{k}^{B}}Q^{*}-Q_{k}^{A}+Q^{*}\}+\alpha w_{k}^{A}\\ &=(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha w_{k}^{A}\\ &\leq(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)(Q_{k}^{A_{U}}-Q^{*})+\alpha w_{k}^{A}\\ &=Q_{k+1}^{A_{U}}-Q^{*},\end{split}$
where the first inequality is due to $\Pi_{Q^{*}}Q^{*}\geq\Pi_{Q_{k}^{B}}Q^{*}$, and the second inequality is due to the hypothesis $Q_{k}^{A_{U}}-Q^{*}\geq Q_{k}^{A}-Q^{*}$. And the second inequality is due to the hypothesis $Q_{k}^{A_{U}}\geq Q_{k}^{A}$ and the fact that the matrix $I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D$ is nonnegative, i.e., all elements are nonnegative by Lemma 1. Therefore, by induction argument, one concludes $Q_{k}^{A_{U}}-Q^{*}\geq Q_{k}^{A}-Q^{*}$ for all $k\geq 0$. The proof of the second inequality follows similar lines. This completes the proof.
C.2 Proof of Proposition 5.2 (Lower comparison system)
Using the dynamic system equation (5.1), we have
$\begin{split}Q_{k+1}^{A}-Q^{*}&=Q_{k}^{A}-Q^{*}+\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{A}-\gamma P\Pi_{Q^{*}}Q^{*}-Q_{k}^{A}+Q^{*}\}+\alpha w_{k}^{A}\\ &=(I-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha D\{\gamma P\Pi_{Q_{k}^{B}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}+\gamma P\Pi_{Q_{k}^{B}}(Q_{k}^{A}-Q_{k}^{B})\}\\ &\quad+\alpha w_{k}^{A}\\ &\geq(I-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha D\{\gamma P\Pi_{Q^{*}}Q_{k}^{B}-\gamma P\Pi_{Q^{*}}Q^{*}+\gamma P\Pi_{Q_{k}^{B}}(Q_{k}^{A}-Q_{k}^{B})\}\\ &\quad+\alpha w_{k}^{A}\\ &=(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})\\ &\quad+\alpha w_{k}^{A}\\ &\geq(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})\\ &\quad+\alpha w_{k}^{A}\\ &=Q_{k+1}^{A_{L}}-Q^{*},\end{split}$
where the first inequality is due to $\Pi_{Q^{B}}Q^{B}\geq\Pi_{Q^{*}}Q^{B}$, and the second inequality is due to the hypothesis $Q_{k}^{A_{L}}-Q^{*}\leq Q_{k}^{A}-Q^{*}$. Therefore, by induction argument, one concludes $Q_{k}^{A_{L}}-Q^{*}\leq Q_{k}^{A}-Q^{*}$ for all $k\geq 0$. And the second inequality is due to the hypothesis $Q_{k}^{{A_{L}}}\leq Q_{k}^{A}$ and the fact that the matrix $I+\alpha\gamma DP\Pi_{Q{*}}-\alpha D$ is nonnegative, i.e., all elements are nonnegative by Lemma 1 The proof of the second inequality follows lines similar to the first proof. This completes the proof.
C.3 Convergence of the lower comparison system
The lower comparison system in (5.3) can be divided into the linear parts with stochastic noises, $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha w_{k}^{A}$ and $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{B_{L}}-Q^{*})+\alpha w_{k}^{B}$, and the external disturbance parts, $\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$ and $\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q_{k}^{A}})(Q_{k}^{A}-Q_{k}^{B})$. As proved in Section B.4, the external disturbances are bounded. Using this fact, one can prove the finite-time error bounds of the linear part with stochastic noise as $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{A_{L}}-Q^{*})+\alpha w_{k}^{A}$ and $(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)(Q_{k}^{B_{L}}-Q^{*})+\alpha w_{k}^{B}$.
Theorem C.1.
For any $k\geq 0$, we have
$\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]$ $\leq\frac{16\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{24\rho^{k-3}k^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}.$ (30)
The same bound holds for $Q_{k}^{B_{L}}-Q^{*}$.
Proof.
First of all, note that (5.3) can be written by
$Q_{k}^{{A_{L}}}-{Q^{*}}=$ ${(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)^{k}}(Q_{0}^{{A_{L}}}-{Q^{*}})+\underbrace{\sum\limits_{j=0}^{k-1}\alpha{{(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)}^{(k-1)-j}}w_{j}^{A}}_{=:(*)}$
$+\underbrace{\alpha\gamma DP\sum\limits_{j=0}^{k-1}{{{(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)}^{(k-1)-j}}}({\Pi_{Q_{j}^{B}}}-{\Pi_{{Q^{*}}}})(Q_{j}^{A}-Q_{j}^{B})}_{=:(**)},$
where ()(*)( ∗ ) reflects the effect of the stochastic noise $w_{j}^{A}$ and ()(**)( ∗ ∗ ) corresponds to the effect of the disturbance $Q_{j}^{A}-Q_{j}^{B}$. Taking the $\infty$-norm on the right-hand side of the above equation leads to
$\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}=$ $\biggl{\lVert}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{k}(Q_{0}^{A_{L}}-Q^{*})$
$\quad+\sum_{j=0}^{k-1}\alpha(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}w_{j}^{A}$
$+\alpha\gamma DP\sum_{j=0}^{k-1}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})(Q_{j}^{A}-Q_{j}^{B})\biggr{\rVert}_{\infty}$
$\leq$ $\biggl{\lVert}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{k}(Q_{0}^{A_{L}}-Q^{*})$
$\quad+\sum_{j=0}^{k-1}\alpha(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}w_{j}^{A}\biggr{\rVert}_{\infty}$
$+\biggl{\lVert}\alpha\gamma DP\sum_{j=0}^{k-1}(I+\alpha\gamma DP\Pi_{Q^{*}}-\alpha D)^{(k-1)-j}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})(Q_{j}^{A}-Q_{j}^{B})\biggr{\rVert}_{\infty}$
$\leq$ $\underbrace{\biggl{\lVert}(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)^{k}(Q_{0}^{{A_{L}}}-{Q^{*}})}_{:=(*)}$
$\underbrace{+\sum\limits_{j=0}^{k-1}\alpha{{(I+\alpha\gamma DP{\Pi_{{Q^{*}}}}-\alpha D)}^{(k-1)-j}}w_{j}^{A}\biggr{\rVert}_{\infty}}_{:=(**)}$
$+\biggl{\lVert}\alpha\gamma DP\biggr{\rVert}_{\infty}\sum_{j=0}^{k-1}\rho^{(k-1)-j}\biggl{\lVert}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})\biggr{\rVert}_{\infty}\biggl{\lVert}(Q_{j}^{A}-Q_{j}^{B})\biggr{\rVert}_{\infty},$
where ()(*)( ∗ ) and ()(**)( ∗ ∗ ) in the second inequality corresponds to the solution of (15) with $x_{k}=Q_{k}^{{A_{L}}}-{Q^{*}}$ and $w_{k}=w_{k}^{A}$, we can apply the bound given in  Corollary A.3. Moreover, applying Lemma 3, one gets
$\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}\leq$ $\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}+\lvert S\times A\rvert^{3/2}\frac{2}{1-\gamma}\rho^{k}$
$\quad+\alpha\gamma d_{\text{max}}\sum_{j=0}^{k-1}\rho^{(k-1)-j}\biggl{\lVert}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})\biggr{\rVert}_{\infty}\biggl{\lVert}(Q_{j}^{A}-Q_{j}^{B})\biggr{\rVert}_{\infty},$
Combining this with (4), we can obtain the following form:
$\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]$ $\leq\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}$
$\quad+\rho^{k-1}\biggl{(}\frac{(k-1)k(2k-1)}{6}\biggl{)}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\biggl{(}\frac{(k-1)k}{2}\biggl{)}\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}$
$\quad+\frac{2\rho^{k}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$ (31)
Then we group some terms of (C.3) as
$\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}\leq 2\biggl{(}\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}\biggl{)}$
Other terms also can be grouped as follows
$\rho^{k-1}\biggl{(}\frac{(k-1)k(2k-1)}{6}\biggl{)}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$+\rho^{k-1}\biggl{(}\frac{(k-1)k}{2}\biggl{)}\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$+\frac{2\rho^{k}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\leq 3\biggl{(}\frac{\rho^{k-3}2k^{3}4\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\biggl{)}$
Then we can get the simplified form as follows
$\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]$ $\leq\frac{16\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{24\rho^{k-3}k^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}$ (32)
C.4 Convergence of the upper comparison system
While the lower comparison system can be analyzed using stochastic linear system characteristic, it is relevantly harder to establish the finite-time error bounds of the upper comparison system because the upper comparison system is a switching system. Therefore, instead of directly finding the finite-time bounds of the upper comparison system, we will use a subtraction system that can be obtained by subtracting the lower comparison system (5.3) from the upper comparison system (5.2) as follows:
$Q_{k+1}^{A_{U}}-Q_{k+1}^{A_{L}}$ $=(I-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP\{\Pi_{Q_{k}^{B}}(Q_{k}^{A_{U}}-Q^{*})-\Pi_{Q^{*}}(Q_{k}^{A_{L}}-Q^{*})\}$
$\quad-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B}),\quad Q_{0}^{A_{U}}-Q_{0}^{A_{L}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$
$Q_{k+1}^{B_{U}}-Q_{k+1}^{B_{L}}$ $=(I-\alpha D)(Q_{k}^{B_{U}}-Q_{k}^{B_{L}})+\alpha\gamma DP\{\Pi_{Q_{k}^{A}}(Q_{k}^{B_{U}}-Q^{*})-\Pi_{Q^{*}}(Q_{k}^{B_{L}}-Q^{*})\}$
$\quad-\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q^{A}})(Q_{k}^{A}-Q_{k}^{B}),\quad Q_{0}^{B_{U}}-Q_{0}^{B_{L}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$ (33)
where the stochastic noises, $w_{k}^{A}$ and $w_{k}^{B}$, are canceled out. If one can prove the stability of the subtraction system, i.e., $Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rightarrow 0$ and $Q_{k}^{B_{U}}-Q_{k}^{B_{L}}\rightarrow 0$ as $k\rightarrow\infty$ then since $Q_{k}^{A_{L}}\rightarrow Q^{*}$ and $Q_{k}^{B_{L}}\rightarrow Q^{*}$ as $k\rightarrow\infty$, one can prove $Q_{k}^{A_{U}}\rightarrow Q^{*}$ and $Q_{k}^{B_{U}}\rightarrow Q^{*}$as $k\rightarrow\infty$ as well. In the following, we prove the finite-time error bound of the subtraction system.
Theorem C.2.
For any $k\geq 0$, we have
$\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]$ $\leq\frac{40\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{1-\gamma}.$ (34)
Proof.
The upper bound of $Q_{k+1}^{A_{U}}-Q_{k+1}^{A_{L}}$ can be presented as following using (C.4)
$Q_{k+1}^{A_{U}}-Q_{k+1}^{A_{L}}$ $=(I-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP\{\Pi_{Q_{k}^{B}}(Q_{k}^{A_{U}}-Q^{*})-\Pi_{Q^{*}}(Q_{k}^{A_{L}}-Q^{*})\}$
$\quad-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$
$=(I-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP\{\Pi_{Q_{k}^{B}}(Q_{k}^{A_{U}}-Q^{*})+\Pi_{Q_{k}^{B}}(Q_{k}^{A_{L}}-Q^{*})$
$\quad-\Pi_{Q_{k}^{B}}(Q_{k}^{A_{L}}-Q^{*})-\Pi_{Q^{*}}(Q_{k}^{A_{L}}-Q^{*})\}-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$
$=(I-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP\Pi_{Q_{k}^{B}}(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})$ (35)
$\quad+\alpha\gamma DP(Q_{k}^{A_{L}}-Q^{*})(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$
$=(I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP(Q_{k}^{A_{L}}-Q^{*})(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})$
$\quad-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$ (36)
Taking the $\infty$-norm on (C.4) and applying the inequality successively result in
$\lVert Q_{k+1}^{A_{U}}-Q_{k+1}^{A_{L}}\rVert_{\infty}$ $\leq\lVert I+\alpha\gamma DP\Pi_{Q_{k}^{B}}-\alpha D\rVert_{\infty}\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}$
$\quad+\biggl{\lVert}\alpha\gamma DP\biggr{\rVert}_{\infty}\sum_{j=0}^{k-1}\rho^{(k-1)-j}\biggl{\lVert}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})\biggr{\rVert}_{\infty}\biggl{\lVert}(Q_{j}^{A_{L}}-Q^{*})\biggr{\rVert}_{\infty}$
$\quad+\biggl{\lVert}\alpha\gamma DP\biggr{\rVert}_{\infty}\sum_{j=0}^{k-1}\rho^{(k-1)-j}\biggl{\lVert}(\Pi_{Q_{j}^{B}}-\Pi_{Q^{*}})\biggr{\rVert}_{\infty}\biggl{\lVert}(Q_{j}^{A}-Q_{j}^{B})\biggr{\rVert}_{\infty}$ (37)
Assuming $Q_{0}^{A_{U}}=Q_{0}^{A_{L}}$ and taking expectation of (C.4) lead to
$\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]$ $\leq\frac{8\gamma^{3}d_{\text{max}}^{3}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+\rho^{k-1}k\frac{2\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}\alpha}{(1-\gamma)}$
$\quad+\rho^{k-1}\frac{(k-1)^{2}k(k-2)}{2}\frac{1}{6}\frac{4\rho^{-3}\alpha^{3}\gamma d_{\text{max}}^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}\frac{1}{2}\frac{k(k-1)(k-2)}{3}$
$\quad+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}$
$\quad+\rho^{k-1}\biggl{(}\frac{(k-1)k(2k-1)}{6}\biggl{)}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\biggl{(}\frac{(k-1)k}{2}\biggl{)}\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$ (38)
We group some terms of (C.4) as follows
$\frac{8\gamma^{3}d_{\text{max}}^{3}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}\leq 5\biggl{(}\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}\biggl{)}$
Also we group other remaining terms as follows
$\rho^{k-1}k\frac{2\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}\alpha}{(1-\gamma)}+\rho^{k-1}\frac{(k-1)^{2}k(k-2)}{2}\frac{1}{6}\frac{4\rho^{-3}\alpha^{3}\gamma d_{\text{max}}^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}\frac{1}{2}\frac{k(k-1)(k-2)}{3}$
$+\rho^{k-1}\biggl{(}\frac{(k-1)k(2k-1)}{6}\biggl{)}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$+\rho^{k-1}\biggl{(}\frac{(k-1)k}{2}\biggl{)}\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\leq 5\biggl{(}\frac{\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}4}{1-\gamma}\biggl{)}$
Then, we can get the following simplified form
$\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]$ $\leq\frac{40\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{1-\gamma}$ (39)
C.5 Proof of Theorem 4.1 (Finite-time error bound of SDQ)
We can use the fact
$\begin{split}\mathbb{E}\left[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}\right]&=\mathbb{E}[\lVert Q_{k}^{A}-Q_{k}^{A_{L}}+Q_{k}^{A_{L}}-Q_{k}^{*}\rVert_{\infty}]\\ &\leq\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]+\mathbb{E}[\lVert Q_{k}^{A}-Q_{k}^{A_{L}}\rVert_{\infty}]\\ &\leq\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]+\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]\end{split}$
The second inequality is due to $Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\geq Q_{k}^{A}-Q_{k}^{A_{L}}.$ This can be inferred from the fact that the lower comparison system and upper comparison system sandwich the original system as $Q_{k}^{L}-Q^{*}\leq Q_{k}-Q^{*}\leq Q_{k}^{U}-Q^{*}$. Then we can rewrite the equation as
$\begin{split}\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]&\leq\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]+\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]\end{split}$
Combining this inequality with (30), (34) yields the following result:
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{16\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{24\rho^{k-3}k^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}$
$\quad+\frac{40\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{1-\gamma}$ (40)
We can group some terms of (C.5) as follows
$\frac{16\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{40\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}\leq 3\biggl{(}\frac{40\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}\biggl{)}$
Other remaining terms can be grouped as follows
$\frac{24\rho^{k-3}k^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{1-\gamma}\leq 2\biggl{(}\frac{24\rho^{k-4}k^{4}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\biggl{)}$
Finally, we can get the finite-time error bound of SDQ
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{120\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\rho^{k-4}k^{4}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
C.6 Proof of Corollary 4.2 (Finite-time error bound of SDQ)
We focus on the term
$\begin{split}k^{4}\rho^{k-4}=\rho^{-4}\rho^{k/2}k^{4}\rho^{k/2}\end{split}$
Let $f(x)=x^{4}\rho^{x/2}$. Solving the first-order optimality condition
$\frac{df(x)}{dx}=\frac{d}{dx}x^{4}\rho^{x/2}=4x^{3}\rho^{x/2}+x^{4}\frac{1}{2}\rho^{x/2}\ln{(\rho)}=0$
we have that its stationary points are $x=\frac{-8}{\ln{(\rho)}}$ and $x=0$. The corresponding function values are
$f\biggl{(}\frac{-8}{\ln{(\rho)}}\biggl{)}=\frac{(-8)^{4}}{(\ln{(\rho)})^{4}}\rho^{\frac{-4}{\ln{(\rho)}}},\quad f(0)=0$
Moreover, solving the second-order optimality condition
$\frac{d^{2}f(x)}{dx^{2}}=\frac{d}{dx}\biggl{(}4x^{3}\rho^{x/2}+x^{4}\frac{1}{2}\rho^{x/2}\ln{(\rho)}\biggl{)}=12x^{2}\rho^{x/2}+4x^{3}\rho^{x/2}\ln{\rho}+\frac{1}{4}x^{4}\rho^{x/2}((\ln(\rho)))^{2},$
we have $f^{\prime\prime}(\frac{-8}{\ln{(\rho)}})<0$ and $f^{\prime\prime}(0)=0$. Therefore, one concludes that $f(\frac{-8}{\ln{(\rho)}})$ is the unique local maximum point. Because the function is continuous and converges to zero as $x\rightarrow+\infty$, it is bounded. This implies that $x=\frac{-8}{\ln{(\rho)}}$ is a global maximum point. Then, we have
$\rho^{(k-4)}k^{4}=\rho^{-4}\rho^{k/2}k^{4}\rho^{k/2}\leq\rho^{-4}{\frac{(-8)^{4}}{(\ln{(\rho)})^{4}}\rho^{\frac{-4}{\ln{(\rho)}}}}\rho^{k/2}$
Combining this bound with (7), one get the bound in (8).